Let f(x)=2x^2-4x+1 and g(x)=3x+2, evaluate the following.
f[g(1)] The answer in the back of the book says 31. How do they get that answer?

Answer by Paul(988)

Its really easy
Plug g(x) into f(x) for every x
And for every x plug in 1

F(g(1))=2[3(1)+2)]-4[3(1)+2]+1
=2(25)-4(5)+1
=50-20+1
=50-19
=31

Paul.